OpenStudy (anonymous):

x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13

OpenStudy (amoodarya):

x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 do you mean system of eq. ?

OpenStudy (amoodarya):

x=3 y=1 z=2 form first eq z=x+5y-6 put it on 2st ,3st eq and solve for x,y